OpenStudy (anonymous):

Challenge: \[f(x)=\sum_{n=1}^{infinty}\sin(\frac{2x}{3^n})\sin(\frac{x}{3^n})\] find f(x) (independent of x) also evaluate the sum of the solutions of the equations f(x)=0 lying in the interval (0,629).

OpenStudy (anonymous):

I didn't make this up I have hundreds of such problems

OpenStudy (anonymous):

What book do you have for these?

OpenStudy (anonymous):

It is more of uncomplied collection

OpenStudy (anonymous):

Any ideas

OpenStudy (anonymous):

sin2x = 2cosx*sinx But it doesn't seem to work

OpenStudy (anonymous):

We shall use transformation formula

OpenStudy (anonymous):

cosc + cosd?

OpenStudy (anonymous):

sinxsin2x=1/2 (cosx-cos3x)

OpenStudy (anonymous):

Oh yeah cosc-cosd, Sorry

OpenStudy (anonymous):

You got it!

OpenStudy (anonymous):

NotSObright you're so bright! I think we have it now

OpenStudy (anonymous):

I know summation formula for cos of angles in AP not GP

OpenStudy (anonymous):

\[f(x) = \cos \frac{x}{3^n} - \cos x, n \to \infty \implies \frac{x}{3^n} \to 0 \implies \cos0 \to 1 \] \[f(x) = 1-\cos{x}\]

OpenStudy (anonymous):

I would like to answer more of such problems

OpenStudy (anonymous):

Thanka man

OpenStudy (anonymous):

f(x)=(1−cosx)/2

OpenStudy (anonymous):

yeah cant forget that 2

OpenStudy (anonymous):

Look at the new thread